Q1 A machine cost $500,000 on January 1, 2006. Its estimated salvage value is $50,000 and its expected life is eight years. Instructions Calculate the depreciation expense (to the nearest dollar) by each of the following methods, showing the figures used.

(a) Double-declining balance for 2007

(b) Sum-of-the-years'-digits for 2007
Answer #1

(a)

Useful life = 8 years

Depreciation rate = (100 / Useful life) X 2

= (100 / 8) X 2

= 25%

Value of machine at the end of 2016 = $500,000 - 25% = $375,000

Depreciation for 2007 = $375,000 X 25% = $93,750

(b)

Useful life is 8 years.

Sum of years = 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 = 36

At the beginning of 2007 the useful life of asset is 7 years.

Depreciation for 2017 = (Cost - Salvage value) X 7/36

= ($500,000 - $50,000) X 7/36

= $87,500 Depreciation for 2007